OpenStudy (anonymous):

An investor invested $1700 in two mutual funds. One fund earned 8% and the other earned 5%. If the total profit was $103, how much was invested in each fund?

OpenStudy (anonymous):

.08x + .05(1700-x) = 103 ( x invested at 8% plus whatever is left of the 1700 for te second fund invested at 5% = profit) .08x + 85 -.05x = 103 (.05 x 1700 = 85) .08x -.05x = 103 -85 (using the order of operations to get the variable) Now, solve for x. Then plus x into (1700-x) to find the amount invested in the second fund.

OpenStudy (anonymous):

*then plug x into (1700-x)...

OpenStudy (anonymous):

Sorry but math is really foriegn to me and I am not sure of how to do that exactly.

OpenStudy (anonymous):

.03x = 18 18/.03 = 600 x=600 1700-600 = 1100 .08(600) + .05(1100) = 48 + 55 = 103 So, the original answer you gave me of 500 and 1200 was incorrect. That would give us a profit of 100, not 103.